OpenStudy (anonymous):

Solve using the elimination method. x+5y=10 and 9x+6y+12. Answer should be in the form (x,y)

OpenStudy (anonymous):

For elimination method, multiply the first equation by 9 and subtract the second one from it as follows: 9 (x + 5y = 10) = 9x + 45y = 90 ----- A 9x + 6y = 12 ----- B Subtracting B from A, we get 39y = 78 so, y = 78/39 so y=2

OpenStudy (anonymous):

Now putting y=2 in x+5y = 10, we get x + 5x2 = 10 x = 10 -10 x = 0 So answer in form (x,y) is (0,2)

OpenStudy (anonymous):

oh. i see thank u

OpenStudy (anonymous):

Got a medal for me??

OpenStudy (anonymous):

always

OpenStudy (anonymous):

Thanks, u r always welcome with yr questions...

OpenStudy (anonymous):

whats the deal with the medals? bragging rights?

OpenStudy (anonymous):

Not really. They're mostly to show whether or not a question has been sufficiently answered.

OpenStudy (anonymous):

You will notice that this question no longer needs a good answer because you gave Harkirat a medal.

OpenStudy (anonymous):

It basically shows that u appreciate the effort of the person who helped u

OpenStudy (anonymous):

As such it won't show up on the list of unanswered questions so that those of us helping people will know we don't need to worry about this thread anymore.

OpenStudy (anonymous):

It's not about effort. It's about sufficiency.

OpenStudy (anonymous):

oh ok. i cant be of much help on this site...my knowledge is in technical areas. but i do appreciate the help i get. :)